# Efficient Biological Data Acquisition through Inference Set Design

### v1

@@ 3 Related Work
In this work, we apply active learning to the problem of hybrid screens for biological data acquisition. Warmuth et al. (2003) were among the firsts to apply active learning to the field of drug discovery, with an acquisition function based on support vector machines for binding affinity predictors. AL has since then been explored as a tool for problems such as virtual screening (Fujiwara et al., 2008), cancer monitoring (Danziger et al., 2009), protein-protein interaction (Mohamed et al., 2010), compound classification (Lang et al., 2016) and chemical space exploration (Smith et al., 2018).
@@ 4 Experiments
In our experiments, we aim at highlighting the key role that inference set design plays into allowing AL agents to reach high levels of performance on the inference set (Section 4.1), validating the robustness of this mechanism on molecular datasets used for drug discovery (Section 4.2) and applying this method to a challenging real-world case of compound library screening on a large scale phenotypic assays (Section 4.3). We focus on comparing our active agent against a random agent, which is the de facto approach used in the industry and often found to be a surprisingly strong baseline even in active learning benchmarks (Yang & Loog, 2018), as well as heuristic-based orderings such as molecular size [^1] The code for all experiments will be made public upon publication of this work. All datasets to reproduce our results are also publicly available, except one proprietary dataset for the results in Figure 10..
4.1 Visualising Inference Set Design
To develop intuition for inference set design we first run experiments on the MNIST dataset. The whole MNIST training set is used as an inference set from which agents can acquire samples. The MNIST test set is split into 50% validation set used for early stopping and 50% test set used for measuring model performance on held-out data inaccessible by agents. In many real-world applications, data used for training ML models contains samples that are inherently difficult, only partially observed, or mislabelled. We investigate these three cases and simulate the challenge of partial observability by removing the lower part of the image (Figure 3b) and simulate labeling noise by shuffling labels for selected digits (see Figure 3c).
Figure 2: Percent of samples with shuffled labels acquired by each agent during active learning experiments.
In these experiments we study the performance of active and random agents. At every active learning step agents acquire a batch of 1000 samples from the inference set, add them to the training set, and retrain the model from scratch. The random agent selects uniformly sampled batches, while the active agent selects samples based on the model’s confidence (see Equation 3). During the active learning experiment we monitor accuracy on the inference set, test set, acquired batch, and the system performance as a whole (see Equation 1). To study the behavior of active agents, we also record what samples are acquired at every active learning step.
@@ 4.1 Visualising Inference Set Design
Figure 3: Performance of active and random agents on variations of the MNIST dataset. a) Original: MNIST dataset with naturally ocurring easy and hard examples. b) Partial Observability: the MNIST dataset where the bottom two thirds of the images have been cropped out. c) Noisy labeling function: MNIST dataset with shuffled labels for 6, 8, and 9 digits.
In all three studied settings in Figure 3, we observe the same phenomenon: the generalisation accuracy measured on the held-out test set quickly saturates for both agents, but the active agent is able to obtain significant improvements on the inference set which are also reflected in individual acquisition batches and on the system’s overall performance. The mechanism at play is the active design of the inference set by the AL agent. We illustrate this for the case of noisy labeling function in Figure 2. By selectively acquiring mislabeled images, the active agent makes predictions on remaining samples in the inference set much easier. Since the majority of samples acquired by an active agent in the first 20 steps are mislabeled, its accuracy on the acquired batch is significantly lower than a random agent’s (Figure 3c, $3$rd row). However, once the process of removing mislabeled samples from the inference set is complete, the active agent’s accuracy on the acquired batch drastically increases and surpasses random agent’s accuracy.
@@
To explore the tradeoff between task difficulty and efficiency gains, we run experiments with shuffled labels for increasingly larger subsets of digits (see Figure 4). The results show that with a higher number of mislabeled samples, the gap between active and random agents’ performance on the inference set increases. The active agent is able to achieve 100% accuracy on the inference set even with shuffled labels for 80% of the data but at a later stage of the experiment. However, a more corrupted target set also necessitates more acquisition batches to prune out the difficult examples from the inference set, incurring smaller experimental budget reductions for highly challenging settings.
We have shown that inference set design that inference set design leads to significantly improved system performance compared to random exploration, even with highly corrupted datasets. In the next section, we show that these observations hold for a wildly different type of data.
Figure 4: Performance of active and random agents on MNIST dataset with different proportions of shuffled labels.
@@ 4.2 Assessing Robustness of the Mechanism
Figure 5: Percent of “hard to predict” molecules acquired by each agent during active learning experiments, as assessed by an independent ensemble.
To demonstrate robustness of the inference set design mechanism, we evaluate our approach on different classification and regression tasks. We use two molecular datasets: Quantum Machine 9 (QM9) (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014) and Molecules3D (Xu et al., 2021). QM9 contains $134k$ small organic molecules and their quantum chemical properties computed with Density Functional Theory (DFT). Molecules3D is a much larger dataset that contains structures and DFT-computed properties of approximately $4$ million molecules.
@@
Figure 6: a) A sample of molecules from the QM9 dataset. b) Performance of active, heuristic, and random agents on HOMO-LUMO gap classification task for QM9 dataset.
In many practical applications exact geometries of screened molecules are unknown as they require computationally expensive DFT calculations. As a first data processing step, we use RDKit library (Landrum et al., 2024) to convert molecular structures into SMILES strings and compute their Extended Connectivity Fingerprints (ECFPs). The SMILES representation provides complete information about molecule’s composition and atomic connectivity, however, it removes all information about 3D atomic positions. Using SMILES representation is a common solution that simplifies generation of candidate molecules for screening but makes property prediction a more challenging task as many properties vary depending on specific 3D conformation of a molecule. We aim to predict total energy for Molecules3D dataset, and HOMO-LUMO gap for both QM9 and Molecules3D.
Datasets are cleaned by removing duplicated SMILES and fingerprints as well as single-atom structures. Since total energies computed by DFT have wide range of large values, we apply additional processing to normalize them (see Appendix C). The final QM9 and Molecule3D datasets contain $133,885$ and $3,453,538$ molecules respectively. Both datasets are split into inference, validation, and test sets with 80%, 5%, 15% fractions. The QM9 HOMO-LUMO gap values are discretized into 2 balanced classes using median as a boundary condition to explore agents’ performance on classification task.
For QM9 dataset, the active agent uses classification uncertainty to select next batch of samples (similar to MNIST experiment). For regression tasks on Molecules3D dataset it uses a query-by-committee approach computing the variance across the predictions of an ensemble. In addition to active and random agents, we evaluate performance of data selection in particular order. For QM9 experiments, we use original sorting of the dataset as acquisition rule, while Molecules3D dataset was sorted from smallest number of atoms in the molecule to largest.
Figure 7: Performance of active, heuristic, and random agents for HOMO-LUMO gap prediction (top) and energy prediction (bottom) on Molecules3D dataset.
The results on QM9 dataset demonstrate that the active agent achieves much higher accuracy on the inference set than the baselines. Furthermore, it reaches near 100% accuracy on both system and inference set after acquiring only 30% of data (see Figure 6). For both HOMO-LUMO gap and total energy regression tasks on Molecules3D dataset active agent achieves approximately 5 times lower MSE compared to random agent or in-order selection by molecule size (see Figure 7). For the QM9 experiments, the accuracy on each acquired batch is used for the stopping criterion.
@@
Similar performance across agents on the test set again suggests that the active agent is able to make gains by pruning out difficult examples from the inference set rather than by generalizing better. To confirm this, we compile a list of “hard to predict” molecules in the QM9 dataset and track what the proportion which is acquired by each agent throughout the experiment. To identify these hard samples, we trained 5 models with different random seeds on a 5-fold cross-validation split of the inference set. We then selected molecules for which at least 4 out 5 models made a mistake. The resulting “hard to predict” subset contains $4,944$ molecules (approximately $5\%$ of the inference set). Figure 5 shows that the active agent acquire molecules from the “hard to predict” subset at a much higher rate than the baselines.
@@
These results shows that inference set design is an effective approach to guide acquisition of molecular compounds. It significantly reduces the cost of acquiring a fixed set of molecules and ensures high system performance. In the next section, we apply this method to real-world biological data.
@@ 4.3 Inference Set Design in the Real World
Modern HTS platforms combined with genetic perturbation techniques facilitate large scale cell microscopy experiments that are designed to capture the effects of biological perturbations (Celik et al., 2022). In these experiments before taking microscopy images, cells undergo perturbation by either a CRISPR/Cas9-mediated gene knock-out or injection of a bioactive molecular compound at given concentration. The obtained cell microscopy images are then processed and passed through a neural network to obtain embeddings that correspond to a specific perturbation. With data from such HTS pipeline it is possible to build “Maps of Biology” that contain organized information about known and new biological relationships (see Figure 8b). Each point on the map corresponds to the cosine similarity between perturbation embeddings and reveals how strongly two perturbations are related. Expanding maps of biology will uncover new biological relationships that in turn will guide the discovery of leads for new medicines. However, even with ultra-HTS platforms, experimentally acquiring microscopy images of all possible cell perturbations is unfeasible. We aim to apply inference set design to this problem with the goal of reducing costs associated with building maps of biology.
@@
Figure 8: a) A microscopy image of perturbed cells from the RxRx3 dataset. b) A sample of biological map from the RXRX3 dataset showing whether each pair of gene-based and compound-based perturbations is pheno-similar (in red) or not (in white).
For our experiments, we start by using the publicly available RXRX3 dataset (Fay et al., 2023) which contains the learned embeddings for $17,063$ CRISPR/Cas9-mediated gene knock-out perturbations, as well as $1,674$ FDA-approved compound perturbations at $8$ concentrations each. We train our models to predict the cosine similarity between the perturbation embeddings of any gene-compound pair, and frame this as a classification problem.
@@
The dataset contains one embedding for each well. Each perturbation type (gene-guide pair or compound-concentration pair) has several replicates across wells, plates and experiments. Each plate also contains unperturbed control cells which are used to keep track of and eliminate a portion of the batch effects (Sypetkowski et al., 2023). These raw embeddings thus need to be aligned and aggregated. We align them by centering and scaling each perturbation embedding to the embeddings of the experiment-level unperturbed control wells. The embeddings are then aggregated through a multi-stage averaging procedure, across wells, plates, experiments and guides (for CRISPR perturbations), which yields an average embedding for each gene-perturbation and each compound-concentration perturbation. For each gene-compound pairs, we select the compound-concentration perturbation that yields the largest cosine-similarity with the target gene. The cosine-similarities are discretized by computing a gene-specific threshold using the upper semi-interquartile range (SIQR) with a step-size of $2.5$. Finally, we remove genes that have activity below 25%. The result is a $151\times 1,674$ matrix of labels encoding either pheno-similarity ($y=1$) or an absence of relationship ($y=0$) between a gene and a compound perturbation (see Figure 8b).
@@
For this prediction task, we treat all gene embeddings as known and focus only on the much larger space of compound acquisition. To featurize compounds we use embeddings of MolGPS model (Sypetkowski et al., 2024). At every active learning step, agents acquire a batch of compounds from the inference set and uncover their relationships with all genes in the dataset (entire row of the matrix). The results on RxRx3 dataset show only minor improvements over random agent (see Appendix E). We believe this is due to high complexity of biological relationships and limited size of the inference set. Indeed, infering cellular similarity between gene and compound perturbations from only a few hundred examples is highly unlikely. To push this analysis further, we run the same experiment but with a much larger proprietary dataset similar to RxRx3 that contains cell perturbation embeddings for 102,855 compound and 5,580 gene perturbations, leading a similarity matrix more than $2,000$ bigger. With inference set design, the active agent is again able to outperform the baselines (see Figure 10). These experiments represent promising results for large scale applications of active learning to hybrid screens in drug discovery programs.
Figure 9: Performance on pheno-similarity classification task for proprietary dataset.
5 Conclusion
@@
We applied this perspective to the problem of hybrid screens in drug discovery, where the goal is to acquire readouts for only a subset of compounds while making accurate predictions for the remaining ones. Our approach makes use of a confidence-based acquisition function and leverages the concept of inference set design, a strategy where the model selects the most challenging examples for labeling, leaving easier cases for prediction. Our empirical results, across image and chemical datasets, as well as a real world biological application, show that this approach leads to consistent and significant improvements.

### v2

@@ 3 Related Work
In this work, we apply active learning to the problem of hybrid screens for biological data acquisition. Warmuth et al. (2003) were among the firsts to apply AL to the field of drug discovery, with an acquisition function based on support vector machines for binding affinity predictors. AL has since then been explored as a tool for problems such as virtual screening (Fujiwara et al., 2008), cancer monitoring (Danziger et al., 2009), protein-protein interaction (Mohamed et al., 2010), compound classification (Lang et al., 2016) and chemical space exploration (Smith et al., 2018).
@@ 4 Experiments
In our experiments, we aim at highlighting the key role that inference set design plays into allowing AL agents to reach high levels of performance on the inference set (Section 4.1), validating the robustness of this mechanism on molecular datasets used for drug discovery (Section 4.2), applying this method to a challenging real-world case of compound library screening on a large scale phenotypic assays (Section 4.3) and empirically validating our theoretical assumption and stopping criterion (Section 4.4). All curves show average metrics across 3 seeds and the shaded area denote standard error around the mean. Vertical lines denote the average stopping time, with min and max intervals (left to right). For all modalities, we use vectorized representations as input and our model $\hat{p}$ is parameterized using a feed-forward deep neural network with residual connections. Additional details regarding data pre-processing, network architectures and hyperparameters are presented in Appendices C and D. The code for all experiments will be made public upon publication of this work. All datasets to reproduce our results are also publicly available, except one proprietary dataset for the results in Figure 8.
@@
Figure 2: Performance of least-confidence (LC) and random acquisition functions on variations of MNIST across 3 seeds. a) Original: MNIST with its naturally ocurring easy and hard examples. b) Partial Observability: MNIST where the bottom two thirds of the images have been cropped out. c) Noisy labeling function: MNIST with shuffled labels for 6, 8, and 9 digits.
Figure 3: Performance of active and random agents on MNIST with increasing number of shuffled labels. As the task becomes harder, the accuracy gap increases, but the stopping time $\tau$ moves back.
4.1 Visualizing Inference Set Design
To develop intuition for inference set design we first run experiments on the MNIST dataset. The whole MNIST training set is used as an inference set from which agents can acquire samples. The MNIST test set is split 50-50 into a validation set used for early stopping and a test set used for measuring model performance on held-out data inaccessible by agents. In many real-world applications, data used for training ML models contains samples that are inherently difficult, only partially observed, or mislabeled. We investigate these three cases and simulate the challenge of partial observability by removing the lower part of the image (Figure 2b) and simulate labeling noise by shuffling labels for selected digits (see Figure 2c).
Figure 4: Percent of samples with shuffled labels acquired by each agent throughout the acquisition steps.
In these experiments we study the performance of an active agent using the least-confidence (LC) acquisition function of Equation 3 and a random agent, selecting samples with uniform probability over the inference set. At every active learning step an agent acquires a batch of 1000 samples from the inference set, adds them to the training set, and retrains the model from scratch. During the active learning experiment we monitor the accuracy on the inference set, test set, acquired batch, and the system accuracy as a whole (see Equation 1). We also record which samples have been acquired at every acquisition steps. In all three studied settings in Figure 2, we observe the same phenomenon: the generalisation accuracy measured on the held-out test set quickly saturates for both agents, but the active agent is able to obtain significant improvements on the inference set which are also reflected in individual acquisition batches and on the system’s overall performance.
@@ 4.1 Visualizing Inference Set Design
The mechanism at play here stems from the active design of the inference set by the active agent. We illustrate this for the case of noisy labeling function in Figure 4. By selectively acquiring the images it tends to mislabel, the active agent makes predictions on remaining samples in the inference set much easier. Since the majority of samples acquired by the active agent in the first 20 steps represent these challenging examples, its accuracy on the acquired batch is significantly lower than a random agent’s (Figure 2c, $3$rd row). However, once the process of removing mislabeled samples from the inference set is complete, the active agent’s accuracy on the acquired batch drastically increases and surpasses the random agent. To further explore the tradeoff between task difficulty and efficiency gains, we run experiments with shuffled labels for increasingly larger subsets of digits (see Figure 3). The results show that with a higher number of such “noisy” samples, the gap between the active and random agents on the inference set increases. The active agent is able to achieve 100% accuracy on the inference set even with shuffled labels for 80% of the data. However, a more corrupted target set also necessitates more acquisition batches to prune out the difficult examples from the inference set, incurring smaller experimental budget reductions for highly challenging settings.
@@ 4.2 Assessing Robustness of the Mechanism
In the previous section, we have shown that inference set design design leads to significantly improved system performance compared to random sampling, even with highly corrupted datasets. In this section, we validate that these observations hold for a wildly different type of data – molecular datasets. We use the Quantum Machine 9 (QM9) (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014). QM9 contains $134k$ small organic molecules and their quantum chemical properties computed with Density Functional Theory (DFT). For inputs, we convert the SMILES strings molecular representations into their Extended Connectivity Fingerprints (ECFPs). We predict the HOMO-LUMO gap framed as a classification task by discetizing the values into two balanced classes using median as a boundary condition to explore agents’ performance on classification task. Our main active agent uses the least-confidence acquisition function. We also explore the use of the BALD acquisition function (Gal et al., 2017). In addition to active and random agents, for both datasets, we also evaluate the performance of heuristic-based acquisition orders (original sorting, smallest to largest molecules, sorted by SA-score, etc.) as well as diversity sampling (selecting molecules with largest Tanimoto distance to the acquired set).
@@
Figure 5: Results on QM9. a) A sample of molecules from the QM9 dataset. b) Performance of active agents (LC and BALD), heuristic orderings, and random sampler on QM9 across 3 seeds.
Figure 6: Percent of “hard” examples acquired by each agent throughout the acquisition loop on QM9.
@@
The results on QM9 demonstrate that both LC and BALD agents achieve much higher accuracy on the inference set than the baselines. Furthermore, they reach near 100% accuracy on both system and inference set after acquiring only 30% of data (see Figure 5). Saturating performance across all agents on the test set suggests once again that the active agent is able to make gains by pruning out difficult examples from the inference set rather than by generalizing better. To confirm this hypothesis, we compiled a list of “hard to predict” molecules by training five models with different random seeds on a 5-fold cross-validation split of the inference set, and flagged the molecules for which at least 4 out 5 models made a mistake. The resulting “hard to predict” subset contains $4,944$ molecules (approximately $5\%$ of the inference set). We then tracked the proportion of these samples that was acquired by each agent in the active learning experiments. Figure 6 shows that the active agents acquire molecules from the “hard to predict” subset at a much higher rate than the baselines.
@@
In addition to the classification results on QM9 presented here, we extend our method to a regression task on the much larger Molecules3D dataset in Appendix E.1 and get to a similar conclusion. Put together, these results on additional datasets spanning both classification and regression tasks and including 3 different active acquisition functions confirm and extend the observations made on MNIST in the previous section. They suggest that inference set design is an effective approach to guide acquisition of molecular compounds, by significantly reducing the cost of acquiring a fixed set of molecules and ensuring high system performance. In the next section, we apply this method to real-world biological data.
@@ 4.3 Inference Set Design in the Real World
Modern HTS platforms combined with genetic perturbation techniques facilitate large scale cell microscopy experiments that are designed to capture the effects of biological perturbations (Celik et al., 2022). In these experiments before taking microscopy images, cells undergo perturbation by either a CRISPR/Cas9-mediated gene knock-out or injection of a bioactive molecular compound at given concentration. The obtained cell microscopy images are then processed and passed through a neural network to obtain embeddings that correspond to a specific perturbation. With data from such HTS pipeline it is possible to build “Maps of Biology” that contain organized information about known and new biological relationships (see Figure 7b). Each point on the map corresponds to the cosine similarity between perturbation embeddings and reveals how strongly two perturbations are related. Expanding maps of biology might uncover new biological relationships that in turn could guide the discovery of leads for new medicines. However, even with ultra-HTS platforms, experimentally acquiring microscopy images of all possible cell perturbations is unfeasible. We aim to apply inference set design to this problem with the goal of reducing costs associated with building maps of biology.
@@
Figure 7: Visualisation of the RxRx3 dataset. a) A microscopy image of perturbed cells. b) A sample of biological map showing whether each pair of gene-based and compound-based perturbations is pheno-similar (in red) or not (in white).
Figure 8: Performance on pheno-similarity classification task on our proprietary phenomics dataset. The least-confidence agent is shown to be the most effective on this task and reaches the targeted system accuracy of 98% at around 80% of the acquisition procedure.
@@
For our experiments, we start by using the publicly available RxRx3 dataset (Fay et al., 2023) which contains the learned embeddings for $17,063$ CRISPR/Cas9-mediated gene knock-out perturbations, as well as $1,674$ FDA-approved compound perturbations at $8$ concentrations each. We train our models to predict the cosine similarity between the perturbation embeddings of any gene-compound pair, and frame this as a classification problem. For each gene-compound pairs, we select the compound-concentration perturbation that yields the largest cosine-similarity with the target gene. The cosine-similarities are discretized by computing a gene-specific threshold using the upper semi-interquartile range (SIQR) with a step-size of $2.5$. Finally, we remove genes that have activity below 25%. The result is a $151\times 1,674$ matrix of labels encoding either pheno-similarity ($y=1$) or an absence of relationship ($y=0$) between a gene and a compound perturbation (see Figure 7b).
@@
For this prediction task, we treat all gene embeddings as known and focus only on the much larger space of compound acquisition. To featurize compounds we use embeddings of MolGPS model (Sypetkowski et al., 2024). At every active learning step, agents acquire a batch of compounds from the inference set and uncover their relationships with all genes in the dataset (entire row of the matrix). The results on RxRx3 dataset show only minor improvements over random agent (see Appendix E). We believe this is due to high complexity of biological relationships and limited size of the inference set. Indeed, inferring cellular similarity between gene and compound perturbations from only a few hundred examples is highly unlikely. To push this analysis further, we rerun the same experiment but on our proprietary phenomics dataset. It is similar to RxRx3 but much larger. It contains cell perturbation embeddings for 102,855 compound and 5,580 gene perturbations, leading a $2,000$ times larger relationship matrix. Using inference set design, the active agent is again able to outperform the baselines on this task, as shown in Figure 8. These experiments represent promising results for large scale applications of active learning to hybrid screens for early drug discovery.
4.4 Empirical validation of weak calibration and stopping criterion
Figure 9: Model calibration evaluation on the inference set in our QM9 experiment for the least-confidence acquisition function. The models tend to be underconfident early in the acquisition loop, and overconfident later on, but remains weakly calibrated (monotonic ordering). At stopping time $t=\tau$, the model is both highly confident and highly accurate on the remaining samples in the inference set.
In this last experimental section, we validate the theoretical claims made in Section 2.2. The proposed stopping criterion assumes weak calibration i.e. that the true accuracy of the model does not need to be exactly equal to its confidence (perfect calibration), but that the ordering is conserved, i.e. that the true accuracy is a monotonically increasing function of the model’s confidence. Figure 9 shows that this assumption is empirically verified in our QM9 experiments and Figure 13 in Appendix F.1 confirms the same observation in our phenomics and MNIST experiments, supporting the claim that weak calibration can be achieved in practice. We also empirically validate (see Figure 14, Appendix F.2) that the proposed stopping criterion making use of the bound presented in Equation 5 does indeed cause the agent to stop acquiring new labels at a system accuracy superior to the targeted threshold (see Equation 4).
5 Discussion
@@
We applied this perspective to the problem of hybrid screens in drug discovery, where the goal is to acquire readouts for only a subset of compounds while making accurate predictions for the remaining ones. Our approach makes use of a confidence-based acquisition function and leverages the concept of inference set design, a strategy where the model selects the most challenging examples for labeling, leaving easier cases for prediction. Our empirical results, across image and chemical datasets, as well as a real world biological application, show that this approach leads to consistent and significant improvements. Moreover, our results highlight that heuristic-based orderings, often used in real-world data acquisition efforts, are highly suboptimal. They can lead to drastically different results across tasks and datasets and often end up harming the performance by introducing unjustified biases in the acquisition ordering. Importantly, this bias then prevents the experimenter from using the performance on the acquisition batch as an indication of the performance on the remaining examples in the inference set. A confidence-based agent also introduces a bias in sample selection, but assuming weak calibration, this bias is directional and still allows the acquisition batch to be used as a lower bound for the performance on the remaining examples (see Lemma 1). It then compensates for the conservativeness of its stopping criterion by designing its own inference set, leading to overall earlier stopping time and greater budget reductions.
An important limitation of our work is the assumption that observed labels are deterministic and constitute perfect accuracy. Many real-world biological experiments however produce noisy observations. Combining the inference set design framework with aleatoric uncertainty modeling to better address cases of weak supervision represents a promising direction for future work.